# FuseFlow: A Fusion-Centric Compilation Framework for Sparse Deep Learning on Streaming Dataflow

### v1

@@ 4. Overview of FuseFlow
Figure 6 summarizes FuseFlow’s compilation flow from PyTorch to a fused, hardware-ready sparse dataflow graph. Models are first lowered from PyTorch (Ansel et al., 2024) to MLIR (Linalg + SparseTensor dialects) using either Torch-MLIR (LLVM, [n. d.]) for dense model components or MPACT (Google, 2021) for sparse model components, with user-specified sparse formats and optional schedules. This process yields a graph of Einsum expressions extended with non-algebraic operators as shown by Figure 6a to Figure 6b. This stage in FuseFlow preserves sparsity semantics from the frontend and provides the knobs (schedules) that guide downstream optimization.
@@
To lower the fused expressions, FuseFlow introduces a new IR called fusion tables (Figure 6d). This representation addresses the complexity of lowering multiple fused expressions to fused dataflow graphs. The fusion table encodes: the fused iteration order based on the partial order graph in its rows (shown in Figure 6d in light pink), the fused expression which is implicit in the tensor order of its columns (shown in off-white), and IR nodes along with named pointers to intermediate streams before they are materialized in their cells (shown in cool gray). Therefore, fusion tables represent the fused tensor computation in a tabular format before code generation, allowing for factored iteration with interleaved input iteration and computation. Because prior work (Hsu et al., 2023, 2025) compiles single kernels, they do not scale to face similar challenges (i.e. requiring references to temporary streams) that FuseFlow address with its fusion table IR.
FuseFlow generates SAM dataflow graphs with ML primitives, which we call SAMML. FuseFlow applies user-guided or autotuned parallelization, sparsity blocking, and dataflow-order selection, and provides a fast heuristic to estimate FLOPs/bytes for early pruning. It then executes in Comal, a cycle-accurate simulator within the open-source DAM simulation framework (Zhang et al., 2024b), or maps to FPGA backends.
Scheduling Language. FuseFlow exposes fusion granularity, dataflow ordering, parallelization, and a performance heuristic through explicit user schedules, enabling the design-space exploration in Section 8. Users specify fusion via Fuse{} regions (i.e. denoted in MLIR with functions), dataflow order via modifying MLIR Linalg affine maps, and other parameters via a command line interface. While this control is essential for optimal performance, future work includes autoscheduling to determine a fusion plan and schedules for common sparse ML patterns.
@@ 5. Cross-Expression Fusion Algorithm
Sparsity-specific challenges arise when multiple expressions each impose their own traversal order over many sparse tensors. Efficient sparse computation requires concordant traversal, a fused iteration order compatible with each operand’s native mode order (storage order) (Zhang et al., 2024a). Traversing a sparse tensor against its storage format (e.g., column-wise over a CSR matrix) is discordant and is often asymptotically worse. Ignoring this critical aspect of sparse computation can lead to incorrect code (Kjolstad et al., 2017) or suboptimal performance due to expensive indirect lookups and tensor reformatting (Zhang et al., 2024a; Kjolstad et al., 2017).
@@
We therefore introduce a fusion algorithm that extends index substitution (de Bruijn, 1972) with ordering constraints maintained in a directed graph called a partial order graph (POG).
@@
2) Build fused Einsum producer-consumer edges: For each tensor, FuseFlow connects producer uses with their consumers, as shown by the arrows in Figure 7, while substituting index variables. This fused Einsum expressions representation is similar to the ideas presented by (Zhou et al., 2022).
@@
By applying the above steps to every expression to be fused, we accumulate a unified index-space representation with all necessary constraints. Throughout this process, the POG ensures that global index ordering remains consistent with all mode orders and user-scheduled dataflow orders encountered. If the graph remains acyclic, FuseFlow performs a topological sort to produce all valid global dataflow orders that respects all constraints. FuseFlow can traverse the fused Einsum representation and emit a single, fully fused Einsum (Figure 7), equivalent to the fused representation in Figure 7 The full algorithm can be found in Section A.5.
@@ 6.1. Fusion Table IR
Figure 9. SAMML graph for the neighborhood subcomputation $T^{nbor}_{ij}=\mathbf{\hat{A}}_{il}X_{lm}\mathbf{\Omega}^{{(2)}}_{mj}$ from GraphSAGE with $i\rightarrow l\rightarrow m\rightarrow j$ order. See Figure 19 in Section A.2 for the full fusion table that generates this graph.
@@
Why use a fusion table?
Conventional compiler representations, such as dependency graphs or value-numbering approaches, represent computations as fixed nodes and edges (Hsu et al., 2023). However, these approaches lack flexibility: once nodes are instantiated, it becomes challenging to reorder or restructure them dynamically without cumbersome graph transformations. In contrast, the fusion table is designed to be a malleable blueprint that the compiler can adjust before final code generation. This brings two key benefits:
@@ 6.2. Code Generation
FuseFlow generates the final dataflow graph by traversing the fusion table top-down, instantiating nodes for coordinate iteration and computation as dictated by the table structure. Starting from the output tensor value cell, FuseFlow recursively expands dependent cells upward, constructing the graph nodes and streams that correspond to the fused loops. Finally, tensor construction nodes (level writers, coordinate droppers) are added to finalize outputs. Figure 9 shows the final dataflow graph generated for a fused GraphSAGE kernel (see Section A.2 for its corresponding fusion table). The result is a hardware-efficient dataflow graph in the factored-iteration style. By design, our fusion table lowering yields a factored (not global) iteration space, interleaving input iteration and computation, as motivated in Section 3 and illustrated in Figure 10 on the right. A direct comparison of the SAM graph with global iteration space and the SAMML graph with factored iteration space is shown in Section A.4. These lowering algorithm implications are further discussed in Section A.4. The full lowering algorithm can be found in Section A.6. Lowering the FuseFlow generated SAMML graph to real hardware follows prior work (Koul et al., 2024), as each node lends itself to VLSI implementations. Therefore, SAMML graphs compose to represent sparse deep learning workloads on dataflow accelerators.
@@ 7. FuseFlow Implementation
We implement FuseFlow within the MLIR compiler framework. We reimplement SAM as an MLIR dialect with a new FuseFlow MLIR compilation path as described in Section 4. Specifically, FuseFlow compiles the MLIR Sparse and Linalg dialects (Bik et al., 2022) to SAM graphs and lets users control fusion and dataflow ordering through a scheduling language as shown in Figure 11. FuseFlow also includes additional optimizations, which were discussed in the context of the SAM IR but not previously present in any SAM compiler (Hsu et al., 2023). These optimizations—such as parallelization, sparsity blocking, dataflow ordering, and an analytical fusion heuristic—are necessary for efficient, large-scale ML. Users can guide these optimizations through a command-line scheduling interface.
PyTorchMLIR: Linalg + SparseTensorFuseFlow CompilerComal SimulatorDataflow AcceleratorScheduling LanguageCollection of Fused CIN RepresentationSAMML IRCross-ExpressionKernel Fusion (Sec. 5)Lowering(Sec. 6)Optimizations(Sec. 7)Refer to captionRefer to captionRefer to caption
Figure 11. System overview of FuseFlow, where blue boxes and arrows denote contributions.
Parallelization. Our compiler applies vectorization and loop unrolling optimizations inspired by SAM (Hsu et al., 2023), concretizing them via stream parallelizer and serializer primitives. Users specify parallelization by selecting index variables and parallelization factors. The compiler partitions tensor coordinates and duplicates compute subgraphs, distributing work across parallel streams and merging results upon completion.
Sparsity Blocking. FuseFlow efficiently targets structured sparsity (e.g., block-sparsity (Zaheer et al., 2020; Dao et al., 2022)) by mapping dense blocks onto the innermost coordinates of tensors. Sparse iteration occurs at outer levels, with dense blocks streamed directly to vectorized ALUs, maintaining sparsity-driven dataflow benefits while enhancing computational density.
Dataflow Ordering. FuseFlow enumerates valid dataflow orders that do not break fusion, enabling users or autotuning frameworks to select schedules to optimize performance. Each dataflow order yields different SAMML graphs and asymptotic efficiencies.
@@ 8. Evaluation
To evaluate the techniques presented in this paper, we use FuseFlow to compile various real-world sparse machine learning applications to our SAMML IR and simulate their end-to-end cycle-accurate performance. We showcase our compiler’s generality by using four different model classes. We also perform ablation studies on various key features of FuseFlow in Section 8.3 to Section 8.6. FuseFlow’s general algorithmic fusion mechanism allows us to explore a vast space of fusion and dataflow schedules, unlocking speedups that were previously unattainable with existing frameworks for sparse workloads on dataflow hardware.
@@ 8.1. Methodology
Benchmark Applications and Datasets. We evaluate FuseFlow on four sparse machine learning model classes across different domains: Sparse Autoencoder (SAE) (Ng et al., 2011) (3 layers), Graph Convolutional Networks (GCN) (Kipf and Welling, 2017) (2 layers), GraphSAGE (Hamilton et al., 2017) (2 layers), and GPT-3 Small (125M parameters) with BigBird attention (Zaheer et al., 2020) (sequence length of 1024). For SAE, we randomly sampled 5 images. Real world datasets used for each model as shown in Section A.3.
@@
Simulator Framework. Our Comal simulator models the architectural behavior of each IR node and tracks cycles based on fully pipelined dataflow graphs, as in SAM (Hsu et al., 2023). It incorporates HBM2 memory simulation via Ramulator 2.0 (Luo et al., 2023), a cycle-accurate DRAM simulator, and provides instrumentation to estimate operations and memory accesses. Comal uses the DAM simulation framework (Zhang et al., 2024b) in Rust 1.87.0 with all simulation functional results verified against a dense PyTorch implementation.
@@ 8.2. Hardware Validation
As in SAM (Hsu et al., 2023), our primary evaluation uses a cycle-accurate simulator. At the time of writing, no existing accelerator broadly supported end-to-end sparse ML. The closest, Onyx (Koul et al., 2024), a coarse-grained reconfigurable array (CGRA) targeting sparse tensor algebra, is insufficient for sparse ML as it lacks support for nonlinear and masking operations. Therefore, we validate simulator fidelity against a post-synthesis RTL simulation of a Xilinx VU9P (AWS F1) design generated from FuseFlow’s SAMML. FuseFlow lowers to Comal for simulation and to Vitis HLS for FPGA, using a minimal, proof-of-concept HLS library to instantiate streaming operators. We select kernels that fit entirely in on-chip BRAM to isolate compute, including partially fused (one-layer) GCN and GraphSAGE, fully fused GCN, and BigBird attention. Concretely, GCN (11 kernels) and GraphSAGE (13 kernels) on KarateClub (Zachary, 1977) and GPT-3 (17 kernels) with sequence length 64. For each model, we normalize kernel cycle counts by the best across both backends and report trend agreement via $R^{2}$ over kernels. We observe a strong agreement of $R^{2}{=}0.991$ in Figure 13. Fairly recently, Chen et al. (2025) introduced Opal, a follow-on CGRA to Onyx with sparse ML support and improved dataflow orderings, which we plan to target as future work.
Figure 13. Latency correlation across kernels across various models. Colors denote models; the dashed line shows parity.

### v2

@@ 4. Overview of FuseFlow
PyTorch, TensorFlow, …MLIR: Linalg + SparseTensorScheduling Language FuseFlow Compiler Collection of Fused CIN Representation SAMML IR FPGA Backend Comal Simulator Dataflow Accelerator Backend HW-aware SAMML IR (koul2025onyx-jssc; onyx)CGRA Dataflow Graph IRCGRA Bitstream (koul2023aha-tecs; koul2025onyx-jssc) HW constrai- nts (hsu2025thesis; koul2025onyx-jssc) Signalelaboration& memory (hsu2025thesis; koul2025onyx-jssc)CGRA (koul2023aha-tecs; koul2025onyx-jssc) Mapping Pipelining & PnR(melchert2024cascade; koul2025onyx-jssc) (koul2023aha-tecs) MPACT (mpact), … Torch-MLIR (torch-mlir), Cross-ExpressionKernel Fusion (Sec. 5)Lowering(Sec. 6)Optimizations(Sec. 7)Refer to captionRefer to captionRefer to caption
Figure 7. System overview of FuseFlow, where blue denotes new contributions.
Figure 6 summarizes FuseFlow’s compilation flow from PyTorch to a fused, hardware-ready sparse dataflow graph. Models are first lowered from PyTorch (pytorch2) to MLIR (Linalg + SparseTensor dialects) using either Torch-MLIR (torch-mlir) for dense model components or MPACT (mpact) for sparse model components, with user-specified sparse formats and optional schedules. Since FuseFlow’s optimizations operate entirely at the MLIR Linalg + SparseTensor dialect level or lower, any frontend, which includes PyTorch through Torch-MLIR (torch-mlir), that lowers to these dialects is supported. This process yields a graph of Einsum expressions extended with non-algebraic operators as shown by Figure 6a to Figure 6b. This stage preserves sparsity semantics from the frontend and provides the knobs (schedules) that guide downstream optimization.
4.1. Supported Sparsity Types
FuseFlow operates on tensors whose sparse structure type is known before compilation, although the data itself does not need to be available until the generated code is executed. The supported sparse data structure types include, in the language of the TACO data structure language (kjolstad2017taco; chou2018), compressed data structures, uncompressed/dense data structures, coordinates, and any combination thereof in higher dimensions (e.g. dense, COO, CSR, DCSR, blocked structures, etc.). This design makes FuseFlow orthogonal to the source of sparsity such as from weights, activations, or inputs.
FuseFlow’s design is sparsity-source agnostic because its fusion algorithm (Section 5) and lowering (Section 6) operate on sparse tensor formats via MLIR’s SparseTensor dialect, which encodes where nonzeros exist. Whether zeros arise from lossless sparsity (e.g., graph adjacency) or lossy sparsity (e.g., magnitude pruning), the resulting compressed format representation is identical. The system’s constraints (Section 5) depend only on tensor mode orders and dataflow dependencies, not on sparsity provenance. Fusion tables (Section 6) organize iteration based on format structure, independent of how that structure was produced.
4.2. Compilation Flow
@@
To lower the fused expressions, FuseFlow introduces a new IR called fusion tables (Figure 6d). This representation addresses the complexity of lowering multiple fused expressions to fused dataflow graphs. The fusion table encodes: the fused iteration order based on the partial order graph in its rows (shown in Figure 6d in light pink), the fused expression which is implicit in the tensor order of its columns (shown in off-white), and IR nodes along with named pointers to intermediate streams before they are materialized in their cells (shown in cool gray). Therefore, fusion tables represent the fused tensor computation in a tabular format before code generation, allowing for factored iteration with interleaved input iteration and computation. Because prior work (hsu2023sam; hsu2025stardust) compiles single kernels, they do not scale to face similar challenges (i.e. requiring references to temporary streams) that FuseFlow address with its fusion table IR.
FuseFlow generates SAM dataflow graphs with ML primitives, which we call SAMML. FuseFlow applies user-guided or autotuned parallelization, sparsity blocking, and dataflow-order selection, and provides a fast heuristic to estimate FLOPs/bytes for early pruning. It then executes in Comal, a cycle-accurate simulator within the open-source DAM simulation framework (dam), or maps to FPGA backends.
Full Compilation Stack. Figure 7 shows the full compilation stack of FuseFlow. Blue components are new contributions of this work and yellow components leverage prior work for the frontend and lower-level compilation paths to hardware. Our contributions focus on algorithms for fusion and other optimizations (PyTorch $\rightarrow$ SAMML graphs), where the challenges lie for performant ML with sparse tensors on dataflow. We describe our compiler implementation and its surrounding software ecosystem in more detail in Section 7.
Scheduling Language. FuseFlow exposes fusion granularity, dataflow ordering, parallelization, and a performance heuristic through explicit user schedules, enabling the design-space exploration in Section 8. Users specify fusion via Fuse{} regions (i.e. denoted in MLIR with functions), dataflow order via modifying Linalg affine maps, and other parameters via a command line interface. While this control is essential for optimal performance, future work includes autoscheduling to determine fusion schedules for common sparse ML patterns.
@@ 5. Cross-Expression Fusion Algorithm
Sparsity-specific challenges arise when multiple expressions each impose their own traversal order over many sparse tensors. Efficient sparse tensor iteration requires concordant traversal, a fused iteration order compatible with each operand’s native mode order (storage order) (zhang2024compilation). Traversing a sparse tensor against its storage format (e.g., column-wise over a CSR matrix) is discordant and is often asymptotically worse. Ignoring this critical aspect of sparse tensor traversal can lead to incorrect code (kjolstad2017taco) or suboptimal performance due to expensive indirect lookups and tensor reformatting (zhang2024compilation; kjolstad2017taco).
@@
We therefore introduce a fusion algorithm that extends index substitution (debruijn1978lambda) with ordering constraints maintained in a directed graph called a partial order graph (POG).
@@
2) Build fused Einsum producer-consumer edges: For each tensor, FuseFlow connects producer uses with their consumers, as shown by the arrows in Figure 8, while substituting index variables. This fused Einsum expressions representation is similar to the ideas presented by (zhou2022react).
@@
By applying the above steps to every expression to be fused, we accumulate a unified index-space representation with all necessary constraints. Throughout this process, the POG ensures that global index ordering remains consistent with all mode orders and user-scheduled dataflow orders encountered. If the graph remains acyclic, FuseFlow performs a topological sort to produce all valid global dataflow orders that respects all constraints. FuseFlow can traverse the fused Einsum representation and emit a single, fully fused Einsum (Figure 8), equivalent to the fused representation in Figure 8 The full algorithm can be found in Section B.4.
@@ 6.1. Fusion Table IR
Figure 10. SAMML graph for the neighborhood subcomputation $T^{nbor}_{ij}=\mathbf{\hat{A}}_{il}X_{lm}\mathbf{\Omega}^{{(2)}}_{mj}$ from GraphSAGE with $i\rightarrow l\rightarrow m\rightarrow j$ order. See Figure 20 in Section B.2 for the full fusion table that generates this graph.
@@
Why fusion tables?
Conventional compiler representations, such as dependency graphs or value-numbering approaches, represent computations as fixed nodes and edges (hsu2023sam). However, these approaches lack flexibility: once nodes are instantiated, it becomes challenging to reorder or restructure them dynamically without cumbersome graph transformations. In contrast, the fusion table is designed to be a malleable blueprint that the compiler can adjust before final code generation. This brings two key benefits:
@@ 6.2. Code Generation
FuseFlow generates the final dataflow graph by traversing the fusion table top-down, instantiating nodes for coordinate iteration and computation as dictated by the table structure. Starting from the output tensor value cell, FuseFlow recursively expands dependent cells upward, constructing the graph nodes and streams that correspond to the fused loops. Finally, tensor construction nodes (level writers, coordinate droppers) are added to finalize outputs. Figure 10 shows the final dataflow graph generated for a fused GraphSAGE kernel (see Section B.2 for its corresponding fusion table). The result is a hardware-efficient dataflow graph in the factored-iteration style. By design, our fusion table lowering yields a factored (not global) iteration space, interleaving input iteration and computation, as motivated in Section 3 and illustrated in Figure 11 on the right. A direct comparison of the SAM graph with global iteration space and the SAMML graph with factored iteration space is shown in Section B.3. These lowering algorithm implications are further discussed in Section B.3. The full lowering algorithm can be found in Section B.5. Lowering FuseFlow’s generated SAMML graph to real hardware follows prior work (onyx), as each node lends itself to VLSI implementations. Therefore, SAMML graphs compose to represent sparse DL on dataflow accelerators.
@@ 7. FuseFlow Implementation
We implement FuseFlow within the MLIR compiler framework. We reimplement SAM as an MLIR dialect with a new FuseFlow MLIR compilation path as described in Section 4. FuseFlow compiles these dialects to SAM graphs and lets users control fusion and dataflow ordering through a scheduling language as shown in Figure 7. FuseFlow also includes additional optimizations, which were discussed in the context of the SAM IR but not previously present in any SAM compiler (hsu2023sam). These optimizations—such as parallelization, sparsity blocking, dataflow ordering, and an analytical fusion heuristic—are necessary for efficient, large-scale ML. Users can guide these optimizations through a command-line scheduling interface.
For lower-level compilation to hardware, we leverage established infrastructure from prior work as shown in Figure 7. This stack handles hardware constraints, signal elaboration, memory allocation, CGRA mapping, and pipelining with place-and-route. Following modern compiler design principles like MLIR, separating high-level transformations from backend-specific lowering enables portability across different dataflow backends (simulator, FPGA, CGRA).
Parallelization. Our compiler applies vectorization and loop unrolling optimizations inspired by SAM (hsu2023sam), concretizing them via stream parallelizer and serializer primitives. Users specify parallelization by selecting index variables and parallelization factors. The compiler partitions tensor coordinates and duplicates compute subgraphs, distributing work across parallel streams and merging results upon completion.
Sparsity Blocking. FuseFlow efficiently targets structured sparsity (e.g., block-sparsity (zaheer2020big; dao2022flashattention)) by mapping dense blocks onto the innermost coordinates of tensors. Sparse iteration occurs at outer levels, with dense blocks streamed directly to vectorized ALUs, maintaining sparsity-driven dataflow benefits while enhancing computational density.
Dataflow Ordering. FuseFlow enumerates valid dataflow orders that do not break fusion, enabling users or autotuning frameworks to select schedules to optimize performance (lacouture_llm_autoscheduling_2025). Each dataflow order yields different SAMML graphs and asymptotic efficiencies.
@@ 8. Evaluation
To evaluate the techniques presented in this paper, we use FuseFlow to compile various real-world sparse machine learning applications to our SAMML IR and simulate their end-to-end cycle-accurate performance. We showcase our compiler’s generality by using four different model classes. We also perform ablation studies on various key features of FuseFlow in Section 8.3 to Section 8.8. FuseFlow’s general algorithmic fusion mechanism allows us to explore a vast space of fusion and dataflow schedules, unlocking speedups that were previously unattainable with existing frameworks for sparse DL on dataflow hardware.
@@ 8.1. Methodology
Benchmark Applications and Datasets. We evaluate FuseFlow on four sparse machine learning model classes across different domains: Sparse Autoencoder (SAE) (ng2011sparse) (3 layers), Graph Convolutional Networks (GCN) (kipf2017semisupervised) (2 layers), GraphSAGE (hamilton2017graphsage) (2 layers), and GPT-3 Small (125M parameters) with BigBird attention (zaheer2020big) (sequence length of 1024). For SAE, we randomly sampled 5 images. Real world datasets (spanning 50%-99.9% and comprising lossless and lossy sparsity sources) were used for each model as shown in Table 2.
Model Dataset MxN Sparsity % Sparsity Source
GCN/GraphSAGE Cora (yang2016revisitingsemisupervisedlearninggraph) 2708x1433 99.7% ZB lossless (in)
GCN/GraphSAGE Cora_ML (bojchevski2018deepgaussianembeddinggraphs) 2995x2879 99.8% ZB lossless (in)
GCN/GraphSAGE DBLP (bojchevski2018deepgaussianembeddinggraphs) 17716x1639 99.6% ZB lossless (in)
GCN/GraphSAGE OGB-Collab (hu2020ogb) 235868x128 99.9% ZB lossless (in)
GCN/GraphSAGE OGB-MAG (hu2020ogb) 1939743x128 99.9% ZB lossless (in)
SAE ImageNet (deng2009imagenet) 224x224 50% ZB lossy (wt)
SAE NIH-CXR (wang2017chestxray) 1024x1024 50% ZB lossy (wt)
SAE LUNA16 (setio2017validation) 512x512 50% ZB lossy (wt)
GPT-3 w/ BigBird IMDB (bojchevski2018deepgaussianembeddinggraphs) 53.9%-86.5%* ZB lossy (mask)
Table 2. Datasets with sparsity levels and types. ZB = zero-based, in = input, wt = weight, mask = masked activation. *Attention mask sparsity.
@@
Simulator Framework. Our Comal simulator models the architectural behavior of each IR node and tracks cycles based on fully pipelined dataflow graphs, as in SAM (hsu2023sam). It incorporates HBM2 memory simulation via Ramulator 2.0 (luo2023ramulator2), a cycle-accurate DRAM simulator, and provides instrumentation to estimate operations and memory accesses. Comal uses the DAM simulation framework (dam) in Rust 1.87.0 with all simulation functional results verified against a dense PyTorch implementation.
@@ 8.2. Hardware Validation
As in SAM (hsu2023sam), our primary evaluation uses a cycle-accurate simulator. At the time of writing, no existing accelerator broadly supported end-to-end sparse ML. The closest, Onyx (onyx), a coarse-grained reconfigurable array (CGRA) targeting sparse tensor algebra, is insufficient for sparse ML as it lacks support for nonlinear and masking operations. Therefore, we validate simulator fidelity against a post-synthesis RTL simulation of a Xilinx VU9P (AWS F1) design generated from FuseFlow’s SAMML. FuseFlow lowers to Comal for simulation and to Vitis HLS for FPGA, using a minimal, proof-of-concept HLS library to instantiate streaming operators. We select kernels that fit entirely in on-chip BRAM to isolate compute, including partially fused (one-layer) GCN and GraphSAGE, fully fused GCN, and BigBird attention. Concretely, GCN (11 kernels) and GraphSAGE (13 kernels) on KarateClub (zachary1977information) and GPT-3 (17 kernels) with sequence length 64. For each model, we normalize kernel cycle counts by the best across both backends and report trend agreement via $R^{2}$ over kernels. We observe a strong agreement of $R^{2}{=}0.991$ in Figure 13. Fairly recently, chen2025opal introduced Opal, a follow-on CGRA to Onyx with sparse ML support and improved dataflow orderings, which we plan to target as future work.
Figure 13. Latency correlation of Comal simulation versus an FPGA across kernels across various models. Colors denote models; the dashed line shows parity.